# multi-AUV; task assigment; kinematic constraints; Dubins path; SOM Neural Network

### v2

@@ 0 Introduction
At present time, there have been a lot of theoretical and practical research results on the task assignment algorithms for multi-AUV system. The task assignment algorithm covers two aspects: 1. task allocation and 2. path planning, which refers to the technology of controlling a group of AUVs according to a certain algorithm and reaching the target along the optimized path under the kinematic constraints. In other words, task allocation ensures optimized task assignment such as Traveling Salesman Problem (TSP); path planning guarantees the feasibility for the agents moving to the assigned tasks. For underwter vehicles, task assignment refers to the assignment of a number of targets to a multi-AUV system, and the cost of the whole multi-AUV system is minimum under the premise that all targets are traversed. Path planning technique refers to the achivement method for any single AUV in the group after global task assignment Huang et al. (2014).
@@
The market mechanism algorithm is one the most common method to solve the multitask assignment operations Oh et al. (2015). Another commonly used task assignment algorithm is ant colony algorithm which mainly studies the autonomous task assignment of multi-robot systems in dynamic environments Cao et al. (2013). These are also other task assignment methods, such as that with intelligent heuristic algorithm handling water flow influence Zhu and Yang (2024), while kinematic path planning not considered for most of them. As for neural networks, literature Liu et al. (2011) further proposed a task allocation algorithm suitable for large-scale multi-robot clusters. The self-organizing neural network algorithm can also be applied to multi-task assignment. The SOM algorithm was proposed by a Finnish scholar, Teuvo Kohonen, in the 1980s Kohonen (1982); Zhao and Guo (2020). Literature Zhu and Yang (2006); Zhu et al. (2013, 2012) applied the self-organizing neural network to the task assignment and distribution of multi-mobile robot systems. path planning. Using the SOM neural network algorithm, dynamic task assignments for multi-AUV systems in two-dimensional ocean environments are implemented. Then, the multi-AUV multi-target allocation strategy of self-organizing map (SOM) neural network is further extended and applied to the three-dimensional marine environment Li and Zhu (2018).
@@
Path planning is one of the intelligent behaviors that AUVs need to possess. The so-called path planning means that in an environment with obstacles, the AUV finds a collision-free path from the initial state to the target state according to certain evaluation criteria Jiang et al. (1999); Leng et al. (2015). Traditional path planning methods mainly include: visual graph method, artificial potential field method, genetic algorithm, fuzzy logic method, etc Khatib (1986); Sugihara (1998); Li et al. (2000). These path planning methods have their own scope of application. However, underwater movements of the AUVs have limitations. If the steering performance by the rudder is limited, with the maximum angle of steering $\theta$, the maximum turning radius should be $R$. For this reason, the maximum turning radius must be smaller than R, which put constrains on the movements. Most of the traditional path planning methods did not consider the kinematic constraints. In another situation, a configurable maximum pitch angle $\phi$ is set to control the behaviour of the climbing of the AUV. In case the necessary pitch angle between two waypoints exceeds this angle, a helical path is introduced to avoid the generation of pitch set-points that might make the vehicle stall. The Dubins path planning method can handle these two kind of situations caused by the vehicle kinematic constraints Cai et al. (2017); Lin and Saripalli (2014). The Dubins method is mainly based on the following theorem. When the direction of motion is known and the turning radius is the smallest, the shortest path from the initial vector to the ending vector is composed of a straight line and a turning arc with the smallest radius. AUV can effectively save energy by using the shortest path in task assignment and path planning. Although the problem of multi-task assignment has been extensively studied, none of the previous studies considered AUV underwater kinematic constraints. The AUV underwater task assignment and path planning in the actual environment cannot ignore the kinematic constraints and the actual existence of the obstacle environment. This paper mainly conducts in-depth research on this. Task assignment and path planning of multiple AUV systems under kinematic constraints in a two-dimensional environment are studied, taking into account obstacles and other interference factors. By combining the Dubins Path algorithm with the improved self-organizing mapping (SOM) neural network algorithm, this paper proposes a new task assignment and path planning algorithm, which effectively solves the multi-task assignment problem of the multi-AUV system in the actual environment.
@@ 2.2 Dubins path planning algorithm for AUV kinematic constrains
Unlike underwater robots such as ROVs, most AUVs and underwater gliders are underactuated with tail thrusters to only moving forward. It is difficult to achieve precise underwater motion control, which poses high requirements for path planning. Fortunately, Dubin’s Car was introduced into the literature by Lester Dubins, a famous mathematician and statistician, in a paper published in 1957. The cars essentially have only 3 controls: “turn left at maximum”, “turn right at maximum”, and “go straight”. All the paths traced out by the Dubin’s car are combinations of these three controls. Dubin’s car is very similar to AUVs’ kinematics. Therefore, we adopt the Dubins method to deal with three-dimensional motion problems.
@@ C Circular arc Na
Dubins curves in above table have been discussed in several papers such as Lin and Saripalli (2014) and Yu Wang et al. (2015), and related calculation methods are provided in former researches, which will not be repeated in this article. Two of the Dubin’s shortest paths don’t use the tangent lines at all. These are the RLR and LRL trajectories, which consist of three tangential, minimum radius turning circles. These trajectories may be used very often in the swarm situation, where a lot of AUVs densely distributed in a certain workspace. If the distance between the agent and the target’s turning circles is less than 4 times the minimum turning radius $r_{min}$, then a CCC curve is valid. For CCC trajectories as an example, we must calculate the location of the third circle, as well as its tangent points to the circles. To be more concrete, the RLR case is shown in Figure 4.
Consider the minimum-radius turning circle to the right of start location to be the circle c1, and the minimum-radius turning circle to the right of the goal location to be the circle c3. The task is now to compute c2, a minimum-radius turning circle tangent to both c1 and c2 plus the points $pt1$ and $pt2$ which are respective points of intersection between c1 and c2, and between c2 and c3. Let p1, p2, p3 be the centers of c1, c2, and c3,respectively. We can form the triangle using these points. Because c2 is tangent to both c1 and c2, we already know the lengths of all three sides.
Segments $\overline{p_{1}p_{2}}$ and $\overline{p_{3}p_{2}}$ have length $2r_{min}$ and segment $\overline{p_{1}p_{3}}$ has length $D$. We are interested in the angle $\theta=\angle p_{3}p_{1}p_{2}$ because that is the angle that the line between c1 and c3 ($\vec{V}_{1}$) must rotate to face the center of c2, which will allow us to calculate p2. Using the law of cosines, we can determine that
$\theta=cos^{-1}\left(\frac{D}{4r_{min}}\right)$ (11)
where $\theta$ in a RLR trajectory represents the amount of rotation that vector $\vec{V}_{1}=p_{3}-p_{1}$ must rotate to point at the center of c2; $D=\sqrt{\left(x_{3}-x_{1}\right)^{2}+\left(y_{3}-y_{1}\right)^{2}}$. However, $\theta$’s value is only valid if $\vec{V}_{1}$ is the same direction as the positive x-axis. Otherwise, the atan2 function will be needed to rotate $\vec{V}_{1}$. For a LRL trajectory, we want to add $\theta$ to this value, but for an RLR trajectory we want to subtract $\theta$ from it to obtain a circle “to the right” of c1. Note that we consider counter-clockwise turns to be positive.
Now that theta represents the absolute amount of rotation, we can compute
$p_{3}=(x_{1}+2r_{min}cos\left(\theta\right),y_{1}+2r_{min}sin\left(\theta\right))$ (12)
@@
where $\vec{V_{2}}=p_{1}-p_{2}$. Next, compute $p_{t1}$ using the new $\vec{V_{2}}$ by
$p_{t1}=p_{3}+\vec{V_{2}}$ (14)
$p_{t2}$ will be computed following a similar procedure.
Figure 4: Computing a RLR trajectory.
@@
After obtaining the dubins path, we modified the task allocation algorithm as follows
@@
With kinematic constraints considered, the computing process increases base on algorithm 1. Due to the limited energy carried by autonomous underwater robots themselves, the rational and effective utilization of energy has become a key constraint for multi AUV systems to complete multiple tasks. Therefore, the load balancing of AUVs is considered in the algorithm. The paper assumes that AUVs carry the same energy, that is, within the same working range, the distances traveled by AUVs exhausting their own energy are the same. Load balancing is determined by the traveling distance limit. At the same time, it is determined whether the distance between the winning AUV and the target task point meets the conditions for forming the Dubins Path. If not, it is set to infinity. Then the input neuron data could be modified in Algorithm 1 to reassign tasks to appropriate AUVs.
@@ 3 Simulation Research
As can be seen in the table above, the distance traveled by AUVs varies depending on whether load balancing is enabled or not. When there is no load balancing, the AUV obtains the minimum total distance based on the operation results of the SOM neural network, but the traveling length of a certain AUV may deviate from the average value. In the case of 4 AUVs for 8 targets, we calculate the mean square deviation of the distance based on the recorded running distance of each AUV, and the calculated formula is $\tilde{L}=\sqrt{\frac{\sum_{i=1}^{n}(L_{i}-\overline{L})^{2}}{n-1}}$. Then the standard deviation of the sample without load balancing can be obtained as 1.812; the standard deviation of the sample with load balancing can be obtained as 0.133. If there is no load balance, the load of a certain AUV may exceed the mean value significantly, which is not feasible in practical applications. Thus, we obtain the multi-task allocation results based on Dubins path planning under load balancing. Under load balancing, due to the fact that the pairing selection between certain AUVs and the target points may not be optimal in distance, the totalpath length will increase and the computation time (tic and toc function) will also increase, which are all within an acceptable range. The simulation results show that the algorithm considering workload balance is effective.
@@
Based on the method discussed in Section II, we extend the application to three-dimensional workspace. Task assigment and Dubins paths in 3D space also face obstacles and load balancing problems. Due to paper length limitations, we will provide an example of simulation here, without considering dynamic disturbances caused by water flow. Simulations were set up with multiple underwater targets deployed randomly in a 3D workspace, as shown in Figure 9(a). Note that in 3D environment the pitch angle $\phi$ is taken into algorithm and restricted to $[-15*PI/180,15*PI/180]$. Task allocation result and trajectoryies are illustrated in Figure 9(b). We have simulated 30 Monte Carlo scenarios with AUVs and targets, and computed the average length of 3D trajectories. The standard deviation value $\tilde{L}$ is relatively small since the proposed algorithms can achieve progressive optimization using load balancing scheme. The comparison results related to energy balance are shown in Table .
@@ 4 Discussion and Future Work
This paper has studied the task assignment and path planning problem for multi-AUV systems with kinematic character considered. The improved SOM neural network method based on workload balance and neighborhood function is adopted to slove the strategy level issues for task allocation. Meanwhile, based on feasible kinematic path planning, combined with the Dubins method for path planning and task reassignment, task allocation and path planning under load balancing are finally achived. 2D and 3D Dubins curves are designed with a set of possible headings and with nonholonomic motion constraints. It is demonstrated that the proposed method has been proven to achieve feasible load balancing for task assignmeng in obstacle environments. In our future work, we will study more practical application and controllers related to dynamic control methods Li et al. (2022), such as Dubins trajectory tracking problem in water flow workspace of multi-AUV systems, as well as collision avoidance of AUVs on the Dubins paths within one group.

### v3

@@ 0 Introduction
At present time, there have been a lot of theoretical and practical research results on the task assignment algorithms for multi-AUV system. The task assignment algorithm covers two aspects: 1. task allocation and 2. path planning, which refers to the technology of controlling a group of AUVs according to a certain algorithm and reaching the target along the optimized path under the kinematic constraints. In other words, task allocation ensures optimized task assignment such as Traveling Salesman Problem (TSP); path planning guarantees the feasibility for the agents moving to the assigned tasks. For underwter vehicles, task assignment refers to the assignment of a number of targets to a multi-AUV system, and the cost of the whole multi-AUV system is minimum under the premise that all targets are traversed. Path planning technique refers to the achivement method for any single AUV in the group after global task assignment Gao et al. (2024).
@@
The market mechanism algorithm is one the most common method to solve the multitask assignment operations Oh et al. (2015). Another commonly used task assignment algorithm is ant colony algorithm which mainly studies the autonomous task assignment of multi-robot systems in dynamic environments Cao et al. (2013). These are also other task assignment methods, such as that with intelligent heuristic algorithm handling water flow influence Zhu and Yang (2024), while kinematic path planning not considered for most of them. As for neural networks, literature Liu et al. (2011) further proposed a task allocation algorithm suitable for large-scale multi-robot clusters. The self-organizing neural network algorithm can also be applied to multi-task assignment. The SOM algorithm was proposed by a Finnish scholar, Teuvo Kohonen, in the 1980s Kohonen (1982); Zhao and Guo (2020). Literature Zhu and Yang (2006); Zhu et al. (2012) applied the self-organizing neural network to the task assignment and distribution of multi-mobile robot systems. path planning. Using the SOM neural network algorithm, dynamic task assignments for multi-AUV systems in two-dimensional ocean environments are implemented. Then, the multi-AUV multi-target allocation strategy of self-organizing map (SOM) neural network is further extended and applied to the three-dimensional marine environment Li and Zhu (2018).
@@
Path planning is one of the intelligent behaviors that AUVs need to possess. The so-called path planning means that in an environment with obstacles, the AUV finds a collision-free path from the initial state to the target state according to certain evaluation criteria Jiang et al. (1999); Leng et al. (2015). Traditional path planning methods mainly include: visual graph method, artificial potential field method, genetic algorithm, fuzzy logic method, etc Khatib (1986); Sugihara (1998); Li et al. (2000). These path planning methods have their own scope of application. However, underwater movements of the AUVs have limitations. If the steering performance by the rudder is limited, with the maximum angle of steering $\theta$, the maximum turning radius should be $R$. For this reason, the minimum turning radius should not be less than R, which put constrains on the movements. Most of the traditional path planning methods did not consider the kinematic constraints. In another situation, a configurable maximum pitch angle $\gamma$ is set to control the behaviour of the climbing of the AUV. In case the necessary pitch angle between two waypoints exceeds this angle, a helical path is introduced to avoid the generation of pitch set-points that might make the vehicle stall. The Dubins path planning method can handle these two kind of situations caused by the vehicle kinematic constraints Cai et al. (2017); Lin and Saripalli (2014). The Dubins method is mainly based on the following theorem. When the direction of motion is known and the turning radius is the smallest, the shortest path from the initial vector to the ending vector is composed of a straight line and a turning arc with the smallest radius. AUV can effectively save energy by using the shortest path in task assignment and path planning. Although the problem of multi-task assignment has been extensively studied, none of the literature mentioned above considered AUV’s underwater kinematic constraints. This issue has certainly been studied before by some scholars from different perspectives, but overall the research is relatively rare. The AUV underwater task assignment and path planning in the actual environment cannot ignore the kinematic constraints and the actual existence of the obstacle environment. This paper mainly conducts in-depth research on this.
As shown in Figure 1, a typical AUV has 1 to 2 thrusters at the tail, relying on rudder and fins to control the moving direction. This structure determines that most AUVs are underactuated. When an AUV moves, it turns horizontally and tilts vertically, i.e., yaw and pitch. The yaw and pitch angles are limited due to mechanical structure and fluid dynamics. In 3D workspace, pitch angle $\gamma\leq\gamma_{max}$, and yaw angle’s derivative or the angular velocity on X-Y plane $\dot{\varphi}$ is also limited, which results the turning radius $r\geq R$, where $R$ is the minimum turning radius under the angular velocity of $max(\dot{\varphi})$.
It should be noted that when the force of the thruster is set very high, the turning radius will also be large at the maximum rudder angle while we are discussing the case of normal thrust.
Considering this, task assignment and path planning of multiple AUV systems under kinematic constraints in a two-dimensional environment are studied in this paper, taking into account static obstacles and other interference factors. The main innovation is to combine the task assignment of SOM neural networks with Dubins path planning under kinematic constraints, making the method more applicable to real-world applications. By combining the Dubins Path algorithm with the improved self-organizing mapping (SOM) neural network algorithm, this paper proposes a new task assignment and path planning algorithm, which effectively solves the multi-task assignment problem of the multi-AUV system in the actual environment. The mathematical model and algorithms are established on MATLAB, and then tested in Python on a Raspberry Pi.
Figure 1: AUV’s kinematic constraint in 3D workspace.
@@ 2.2 Dubins path planning algorithm for AUV kinematic constrains
Unlike underwater robots such as ROVs, most torpedo like AUVs and underwater gliders are underactuated with tail thrusters to only moving forward, steering depend on fins and rudders. It is difficult to achieve precise underwater motion control, which poses high requirements for path planning. Fortunately, Dubin’s Car was introduced into the literature by Lester Dubins, a famous mathematician and statistician, in a paper published in 1957. The cars essentially have only 3 controls: “turn left at maximum”, “turn right at maximum”, and “go straight”. All the paths traced out by the Dubin’s car are combinations of these three controls. Dubin’s car is very similar to AUVs’ kinematics. Having precise motion control to solve dynamic problems is also very useful, which is another quite challenging field being studied and will be involved in our future work. Here, we adopt the Dubins method to deal with three-dimensional motion problems.
@@ C Circular arc Na
Dubins curves in above table have been discussed in several papers such as Lin and Saripalli (2014) and Yu Wang et al. (2015), and related calculation methods are provided in former researches, which will not be repeated in this article. Two of the Dubin’s shortest paths don’t use the tangent lines. These are the RLR and LRL trajectories, which consist of three tangential, minimum radius turning circles Giese (2012–2023); Shkel and Lumelsky (2001). These trajectories may be used very often in the swarm situation, where a lot of AUVs densely distributed in a certain 3D workspace. If the distance between the agent and the target’s turning circles is less than 4 times the minimum turning radius $r_{min}$, then a CCC curve is valid. For CCC trajectories as an example, we must calculate the location of the third circle, as well as its tangent points to the circles. To be more concrete, the RLR case is shown in Figure 5.
Consider the minimum-radius turning circle to the right of start location to be the circle c1, and the minimum-radius turning circle to the right of the goal location to be the circle c3. The task is now to compute c2, a circle tangent to both c1 and c2 plus the points $p_{t}1$ and $p_{t}2$ which are respective points of intersection between the 3 circles. Let p1, p2, p3 be the centers of circle c1, c2, and c3,respectively. The triangle is formed using these points. Because c2 is tangent to both c1 and c2, the lengths of all three sides are known.
[…]
@@
[…]
Figure 5: Computing a RLR Dubins trajectory.
@@
After obtaining the dubins path, we modified the task allocation algorithm as following Figure 6. With kinematic constraints considered, the computing process increases base on algorithm 1. Due to the limited energy carried by autonomous underwater robots themselves, the rational and effective utilization of energy has become a key constraint for multi AUV systems to complete multiple tasks. Therefore, the load balancing of AUVs is considered in the algorithm. The paper assumes that AUVs carry the same energy, that is, within the same working range, the distances traveled by AUVs exhausting their own energy are the same. Load balancing is determined by the traveling distance limit. At the same time, it is determined whether the distance between the winning AUV and the target task point meets the conditions for forming the Dubins Path. If not, it is set to infinity. Then the input neuron data could be modified in Algorithm 1 to reassign tasks to appropriate AUVs.
@@ 3 Simulation Research
As can be seen in the table above, the distance traveled by AUVs varies depending on whether load balancing is enabled or not. The algorithms were firstly implemented with MATLAB R2016a software on the hardware platform of HP Computer with Intel Core i7 9700 CPU and 16G DDR4 2400 memory. Simulation run time consumption was recorded using the Tic and Toc function. When there is no load balancing, the AUV obtains the minimum total distance based on the operation results of the SOM neural network, but the traveling length of a certain AUV may deviate from the average value. In the case of 4 AUVs for 8 targets, we calculate the mean square deviation of the distance based on the recorded running distance of each AUV, and the calculated formula is $\tilde{L}=\sqrt{\frac{\sum_{i=1}^{n}(L_{i}-\overline{L})^{2}}{n-1}}$. Then the standard deviation of the sample without load balancing can be obtained as 1.812; the standard deviation of the sample with load balancing can be obtained as 0.133. If there is no load balance, the load of a certain AUV may exceed the mean value significantly, which is not feasible in practical applications. Thus, we obtain the multi-task allocation results based on Dubins path planning under load balancing. Under load balancing, due to the fact that the pairing selection between certain AUVs and the target points may not be optimal in distance, the total path length will increase and the computation time will also increase, which are all within an acceptable range on the level of tens of milliseconds. The simulation results show that the algorithm considering workload balance is effective.
Meanwhile, we implemented the algorithm through Python programs, with package NumNy. On the PC, the simulation time consumption with Python is approximately 5-10 times that of Matlab. In order to get closer to practical applications, we run Python simulations on the embedded device, an official Raspberry Pi 4B with 4G memory. The running times are shown in the table, which basically take a few seconds to complete the core algorithm, without drawing the graphics. Considering the rapid development of ARM computing power, it can be said feasible to run this algorithm on embedded systems. It should be noted that running on the Pi is only an experiment for the core program, not in an actual operation. The field applications may require an omniscient perspective and a known global map (or SLAM), which may involve various additional hardware.
@@
Based on the method discussed in Section II, we extend the application to three-dimensional workspace. Task assigment and Dubins paths in 3D space also face obstacles and load balancing problems. Due to paper length limitations, we will provide an example of simulation here, without considering dynamic disturbances caused by water flow. Simulations were set up with multiple underwater targets deployed randomly in a 3D workspace, as shown in Figure 10(a). Note that in 3D environment the pitch angle $\gamma$ is taken into algorithm and restricted to $[-15*PI/180,15*PI/180]$. Task allocation result and trajectoryies are illustrated in Figure 10(b). We have simulated 30 Monte Carlo scenarios with AUVs and targets, and computed the average length of 3D trajectories. The standard deviation value $\tilde{L}$ is relatively small since the proposed algorithms can achieve progressive optimization using load balancing scheme. The comparison results related to energy balance are shown in Table 3. We also implemented some simulations on PC (MATLAB) and Pi (Python). The average running time within an acceptable range has been recorded in the table.
@@
Figure 10: Task assignment in 3D environment with workload balancing.
Table 3: Task assignment and path planning results in 3D workspace with and without workload balancing.
@@ 4 Discussion and Future Work
This paper has studied the task assignment and path planning problem for multi-AUV systems with kinematic character considered. The improved SOM neural network method based on workload balance and neighborhood function is adopted to slove the strategy level issues for task allocation. Meanwhile, based on feasible kinematic path planning, combined with the Dubins method for path planning and task reassignment, task allocation and path planning under load balancing are finally achived. 2D and 3D Dubins curves are designed with a set of possible headings and with nonholonomic motion constraints. It is demonstrated that the proposed method has been proven to achieve feasible load balancing for task assignmeng in obstacle environments. In future work, firstly, we will rewrite the algorithm in C++, hoping to achieve faster speed on embedded systems. Secondly, we will study more practical application and controllers related to dynamic control methods Li et al. (2022), such as Dubins trajectory tracking problem in water flow workspace of multi-AUV systems, as well as inter group collision avoidance of AUVs on the Dubins paths, and the moving obstacle avoidance problem.